### v1

# A Survey of Deep Graph Learning under Distribution Shifts: from Graph Out-of-Distribution Generalization to Adaptation

Abstract
Distribution shifts on graphs – the discrepancies in data distribution between training and employing a graph machine learning model – are ubiquitous and often unavoidable in real-world scenarios. These shifts may severely deteriorate model performance, posing significant challenges for reliable graph machine learning. Consequently, there has been a surge in research on graph machine learning under distribution shifts, aiming to train models to achieve satisfactory performance on out-of-distribution (OOD) test data. In our survey, we provide an up-to-date and forward-looking review of deep graph learning under distribution shifts. Specifically, we cover three primary scenarios: graph OOD generalization, training-time graph OOD adaptation, and test-time graph OOD adaptation. We begin by formally formulating the problems and discussing various types of distribution shifts that can affect graph learning, such as covariate shifts and concept shifts. To provide a better understanding of the literature, we systematically categorize the existing models based on our proposed taxonomy and investigate the adopted techniques behind. We also summarize commonly used datasets in this research area to facilitate further investigation. Finally, we point out promising research directions and the corresponding challenges to encourage further study in this vital domain. Additionally, we provide a continuously updated reading list at https://github.com/kaize0409/Awesome-Graph-OOD.
Index Terms:
Graph Learning, Distribution Shifts, Graph Out-of-Distribution Generalization, Graph Out-of-Distribution Adaptation.
1 Introduction
Driven by the prevalence of graph-structured data in numerous real-world scenarios, growing attention has been paid to graph machine learning, which effectively captures the relationships and dependencies among entities within graphs. In particular, Graph Neural Networks (GNNs) have emerged as powerful tools for learning representations on graphs through message-passing [1, 2, 3], and they have demonstrated remarkable success across diverse applications, such as social networks, physics problems, and traffic networks [4, 5, 6].
While graph machine learning has achieved notable success, most of the existing efforts assume that test data follows the same distribution as training data, which is often invalid in the wild. When confronted with Out-Of-Distribution (OOD) samples, the performance of graph machine learning methods may substantially degrade, limiting their efficacy in high-stake graph applications such as finance and healthcare [7]. Although numerous transfer learning methods have been proposed to address distribution shifts for Euclidean data [8, 9, 10], their direct application to graph data is challenging. This is due to the interconnected nature of entities on graphs, which violates the independent and identically distributed (IID) assumption inherent in traditional transfer learning methods. Moreover, the various types of graph shifts introduce new challenges. These shifts occur across different modalities including features, structures, and labels, and can manifest in various forms such as variations in graph sizes, subgraph densities, and homophily [11]. Given these obstacles, increasing research efforts have been dedicated to improving the reliability of graph machine learning against distribution shifts, concentrating on three main scenarios: graph OOD generalization  [7, 11], training-time graph OOD adaptation [12, 13], and test-time graph OOD adaptation [14, 15].
The primary distinction between graph OOD generalization and adaptation methods lies in their assumptions regarding the availability of target data. Graph OOD generalization methods typically assume the unavailability of target data during model training and aim to enhance the model’s generalization performance on any potential unseen test distribution. In contrast, both training-time and test-time adaptation methods assume the availability of target data and aim to improve model performance on this specific target. However, they differ in their assumptions about the source data and in how they utilize knowledge of the source distribution. Training-time adaptation assumes that both the source and target graphs are available simultaneously, allowing model adaptation to start from scratch during the training process. On the other hand, test-time adaptation typically assumes access to a model pre-trained on the source graph, rather than the source graph itself, and begins adapting the model to the target data from this pre-trained state. Although graph OOD generalization, training-time OOD adaptation, and test-time OOD adaptation are closely related, there is currently no unified framework that comprehensively discusses deep graph learning under distribution shifts across all three scenarios.
With recent progress on graph OOD learning, an up-to-date and forward-looking review of this field is urgently needed. In this survey, we provide, to the best of our knowledge, the first unified and systematic review of the literature on deep graph learning under distribution shifts. We start by formally formulating the problems and discussing different types of graph distribution shifts in graph machine learning. Next, our new taxonomy is proposed, classifying existing methods into three categories based on the model learning scenario: (1) graph OOD generalization, where the generalizability is enhanced through strategic design of the model when training on source data, (2) training-time graph OOD adaptation, where the adaptation is performed when jointly training the model based on both source and target data [16, 17], and (3) test-time graph OOD adaptation, where the adaptation happens when adjusting a pre-trained source model to the target data [18, 19].
To deepen our understanding of these approaches, we further classify existing methods within each of the three categories into model-centric and data-centric strategies. Model-centric approaches focus on the learning process or the architecture of the graph model itself, enhancing the model’s inherent ability to generalize or adapt to distribution shifts by refining its structure, training objectives, or learning mechanisms. In contrast, data-centric approaches emphasize the manipulation of input graphs, improving model performance by addressing the data directly, either through preprocessing techniques or data augmentation strategies. Within each subline of research, we elaborate on the detailed techniques for enhancing the generalizability or adaptability under distribution shifts on graphs. Additionally, we provide a summary of the datasets utilized in these studies, highlighting their characteristics and relevance to the challenges posed by distribution shifts. Based on the current progress on graph OOD learning, at the end we also point out several promising research directions in this evolving field.
Differences between this survey and existing ones. Despite the urgent need for an overview of graph learning under distribution shifts, existing surveys have primarily focused on subfields within this area, rather than providing a comprehensive overview from multiple scenarios. Until now, there have been several surveys in related areas, but exhibiting distinct focuses, including graph OOD generalization [7, 20], graph domain adaptation [21, 22], trustworthy graph learning [23] related to distribution shifts. Our work distinguishes itself from existing ones in the following aspects: (1) Main focus. Our survey centers on the challenges and solutions for graph learning under distribution shifts, while [23] analyzes OOD issues from a trustworthy perspective and does not delve into the methodological aspects. Conversely, [20] examines graph machine learning from a causal perspective, which is narrower than our broad examination. (2) Taxonomy. We provide a comprehensive categorization of existing methods and summarize them, whereas related work, such as [24], lacks such summaries. Other surveys like [21] and [22] primarily focus on domain adaptation without addressing the broader scope of graph OOD learning. Additionally, we provide coverage of the most recent advancements and discussions in this field.
Survey structure. The general organization of this survey is presented as follows: Section 2 introduces the notations and preliminaries. Sections 3, 4 and 5 review graph OOD generalization, training-time graph OOD adaptation, and test-time graph OOD adaptation, respectively. Each section discusses model-centric and data-centric approaches within its scenario, further detailing the techniques associated with each category. Furthermore, Section 6 provides a comprehensive summary of the datasets used in the literature, highlighting popular graph datasets for evaluation and their relevance to the challenges posed by distribution shifts. Section 7 explores promising future research directions and the associated challenges. Finally, Section 8 presents the conclusion of this survey.
2 Preliminaries
2.1 Problem Definition
Figure 1: An illustration of graph OOD generalization, training-time graph OOD adaptation, and test-time graph OOD adaptation. The source graph can be single or multiple.
Let $\mathcal{V}$ denote the node set with $N$ nodes in a graph $\mathcal{G}=(\textbf{A},\textbf{X})$, where $\textbf{A}\in\mathbb{R}^{N\times N}$ is the adjacency matrix, with $a_{uv}$ representing the connection between nodes $i$ and $j$. The node feature matrix $\textbf{X}\in\mathbb{R}^{N\times F}$ contains $F$-dimensional feature vectors for each node $v\in\mathcal{V}$. Additionally, $\textbf{Y}\in\mathbb{R}^{N\times K}$ represents the node label matrix, where each label $y_{v}$ corresponds to one of the $K$ classes.
In this section, we take node classification as an exemplar to define node-level tasks of graph learning under distribution shifts. We consider three distinct problem scenarios (i.e., Graph OOD Generalization, Training-time Graph OOD Adaptation, and Test-time Graph OOD Adaptation) and the illustration can be found in Figure 1. Note that these definitions can be readily extended to encompass edge-level and graph-level learning tasks. Specifically, let $\varphi_{\theta}$ denote a graph model characterized by parameters $\theta$. Generally, the deep graph model $\varphi$ can be decomposed as $g\circ f$, where $g(\cdot):(\textbf{A},\textbf{X})\rightarrow\textbf{H}$ is a graph encoder mapping the graph inputs to latent representations H, and $f(\cdot)$ is a classifier in the latent space.
Problem Definition 1
Graph OOD Generalization: In this scenario, we assume that distribution shifts exist between source and target distributions $\mathcal{P}_{S}\neq\mathcal{P}_{T}$, and that $\mathcal{P}_{T}$ is not observable during model training. The objective of node-level graph OOD generalization is to learn an optimal model $\varphi_{{\theta}^{*}}$ using the nodes and labels from the source graph $(\mathcal{G}_{S},\textbf{Y}_{S})$, such that it generalizes well to the unseen target graph $\mathcal{G_{T}}$.
$\varphi_{{\theta}^{*}}=\arg\min_{\theta}\mathbb{E}_{\mathcal{P}_{S}}[l(\varphi_{\theta}(\mathcal{G}_{S}),\textbf{Y}_{S})],$ (1)
where $l(\cdot)$ is the loss function that measures the discrepancy between the model’s predictions $\varphi_{\theta}(\mathcal{G}_{S})$ and the ground-truth node labels $\textbf{Y}_{S}$. The learned model $\varphi_{{\theta}^{*}}$ is then used to infer node labels $\hat{\textbf{Y}}_{T}=\varphi_{{\theta}^{*}}(\mathcal{G}_{T})$ for the target graph.
Problem Definition 2
Training-time Graph OOD Adaptation: This scenario involves a source graph $\mathcal{G}_{S}$ with corresponding node labels $\textbf{Y}_{S}$, and a target graph $\mathcal{G}_{T}$ with node labels $\textbf{Y}_{T}$, where $\textbf{Y}_{T}$ may be partially labeled or entirely unlabeled (i.e., $\textbf{Y}_{T}=\varnothing$). Under the assumption that there exist distribution shifts between the source and target distributions, the objective of node-level training-time graph OOD adaptation is to learn an optimal model $\varphi_{{\theta}^{*}}$ that minimizes the loss on the target graph by leveraging both the source and target graphs:
$\varphi_{{\theta}^{*}}=\arg\min_{\theta}\mathbb{E}_{\mathcal{P}_{S},\mathcal{P}_{T}}[l(\varphi_{\theta}(\mathcal{G}_{S},\mathcal{G}_{T}),\textbf{Y}_{S},\textbf{Y}_{T})].$ (2)
If $\textbf{Y}_{T}=\varnothing$, indicating that none of the nodes in the target graph are labeled, the problem is classified as unsupervised domain adaptation. The learning objective $l(\cdot)$ exploits source domain labels but would not rely on target domain labels. Otherwise, when a portion of the nodes in the target graph are labeled, the problem is considered semi-supervised, and $l(\cdot)$ can utilize both the source domain labels and the partial target domain labels to optimize model performance.
Problem Definition 3
Test-time Graph OOD Adaptation: In this scenario, a pre-trained model $\varphi_{\theta^{\prime}}$ and a target graph $\mathcal{G_{T}}$ with node labels $\textbf{Y}_{T}$ (which may be partially labeled or entirely unlabeled) are given. The goal is to adapt the pre-trained model during the test-time training phase in response to distribution shifts in the target domain. This adaptation is performed without access to source domain data during this phase. Given the distribution shifts between the source and target domains, the objective is to adjust the pre-trained model $\varphi_{\theta^{\prime}}$ at test time to optimize its performance on the target graph:
$\varphi_{{\theta}^{*}}=\arg\min_{\theta}\mathbb{E}_{\mathcal{P}_{T}}[l(\varphi_{\theta^{\prime},\theta}(\mathcal{G}_{T}),\textbf{Y}_{T})].$ (3)
Here, $\theta^{\prime}$ represents the parameters from the pre-trained model, and $\theta$ represents the parameters introduced for test-time adaptation.
Figure 2: Overview of our proposed taxonomy. Within each of the three scenarios, existing methods can be further categorized into model-centric and data-centric strategies.
Graph Distribution Shift Types. In traditional machine learning, several studies have discussed and defined various types of distribution shifts [25, 26], of which the most widely-used concepts are covariate shifts (i.e., shifts in $\mathcal{P}(\textbf{X})$) and concept shifts (i.e., shifts in $\mathcal{P}(\textbf{X}|\textbf{Y})$ or $\mathcal{P}(\textbf{Y}|\textbf{X})$). These concepts can naturally be extended to graph setting by replacing feature inputs X with the graph inputs $\mathcal{G}=(\textbf{A},\textbf{X})$.
  • Covariate Shifts. Covariate shifts on graphs emphasize the changes in graph inputs $\mathcal{P}(\textbf{A},\textbf{X})$. This can be further decomposed and interpreted as structure covariate shift, size covariate shift, and feature covariate shift [7]. Specifically, size covariate shift refers to changes in the graph size distribution, which can be captured as shifts in $\mathcal{P}(\textbf{A},\textbf{X},N)$, where $N$ denotes the graph size [27].
  • Concept Shifts. Concept shifts on graphs highlight the shifts in the relationship between graph inputs and labels $\mathcal{P}(\textbf{Y}|\textbf{A},\textbf{X})$ or $\mathcal{P}(\textbf{A},\textbf{X}|\textbf{Y})$. The concept shifts can be further decomposed to reveal more specific graph distribution shift types, such as feature concept shift $\mathcal{P}(\textbf{X}|\textbf{Y})$ and structure concept shift $\mathcal{P}(\textbf{A}|\textbf{Y})$ [5].
The usage of these concepts often extends beyond the input space to the latent representation space H, with covariate shifts describing the distribution shifts in latent representations $\mathcal{P}(\textbf{H})$, and concept shifts describing the changes in $\mathcal{P}(\textbf{H}|\textbf{Y})$ or $\mathcal{P}(\textbf{Y}|\textbf{H})$.
2.2 Discussion on Related Topics
Several topics are closely related to graph OOD learning, including: graph transfer learning, graph domain adaptation, graph OOD detection, and graph continual learning. These topics, sharing similar goals with graph OOD learning but exhibiting nuanced differences, are discussed as follows.
Graph Transfer Learning. Traditionally, transfer learning is a more wide-ranging concept and involves transferring knowledge across distribution changes and diverse tasks. Compared to the focus of this survey – graph learning under distribution shifts, graph transfer learning does not exclusively focus on addressing distribution shifts, but also includes leveraging knowledge acquired from one task to enhance performance in another context.
Graph Domain Adaptation. Following traditional domain adaptation, graph domain adaptation methods [21, 22] typically rely on the covariate shift assumption about an invariant relationship between graph inputs and labels, represented as $\mathcal{P}_{S}(\textbf{Y}|\textbf{A},\textbf{X})=\mathcal{P}_{T}(\textbf{Y}|\textbf{A},\textbf{X})$. The objective is then to address distribution shifts in the input graph space across domains $\mathcal{P}_{S}(\textbf{A},\textbf{X})\neq\mathcal{P}_{T}(\textbf{A},\textbf{X})$. In contrast to graph domain adaptation, graph OOD learning is more general and comprehensive, involving distribution shifts that go beyond the covariate shift assumption.
Graph OOD Detection. Graph OOD detection aims to empower models to recognize samples that significantly deviate from the distribution of the training data. In particular, the open-set problem, wherein unseen classes emerge in the test data, can be regarded as a special case of the OOD detection problem by defining the OOD nodes as nodes with labels unseen from the training set [28]. Tackling distribution discrepancy in the open-set setting typically involves the dual objective of performing OOD detection to identify uncertain nodes while also adapting the model to the remaining non-OOD nodes [29, 30].
Graph Continual Learning. Graph continual learning involves adapting to a continuous stream of evolving graph data while retaining previously acquired knowledge [31, 32]. This process focuses on the model’s ability to incrementally learn and update itself as new data arrives, without forgetting past information. In contrast, graph OOD learning is primarily concerned with managing distribution shifts in static data and is focused on specific scenarios where such shifts occur.
2.3 Proposed Taxonomy
As seen from the problem definitions, graph OOD generalization, training-time graph OOD adaptation, and test-time graph OOD adaptation, significantly differ in their model learning scenarios. In particular, graph OOD generalization methods assume the inaccessibility of target data and involve only source data; training-time graph OOD adaptation methods involve both source and target data and start model training from scratch, while test-time graph OOD adaptation methods assume the unavailability of source data and aim to adapt a pre-trained source model to the target. Accordingly, in the following three sections, we firstly categorize existing methods into graph OOD generalization, training-time graph OOD adaptation, and test-time graph OOD adaptation. Within each section, we further classify methods into model-centric and data-centric approaches. Model-centric approaches center on the learning process or the design of the graph model, while data-centric approaches emphasize the manipulation of input graphs, such as transforming graph structures or features. Our taxonomy is illstrated in Figure 2.
3 Graph OOD Generalization
In the field of graph machine learning against distribution shifts, the scenario of graph OOD generalization problems has attracted increasing attention in recent years [33, 34, 35, 36]. The setting of graph OOD generalization assumes that the model has no access to the target domain during training, which is notably more challenging than the domain adaptation setting, where the target domain is known in advance [5, 18]. Therefore, it is particularly crucial to enhance the overall generalizability of the model, such that it can handle samples from any unseen domain [37]. In this section, we provide a comprehensive review of the most recent advances in graph OOD generalization, focusing on the model-centric and data-centric approaches. A summary of existing methods is presented in Table I.
3.1 Model-Centric Approaches
In this section, we introduce model-centric approaches for graph OOD generalization, which focus on the strategies in model learning to improve generalizability. Based on specific designs and objectives, these approaches could be further categorized into invariant representation learning, causality-based learning and graph self-supervised learning.
Invariant Representation Learning. Invariant representation learning methods for graph OOD generalization aim to extract invariant information that is robust to distribution shifts across different environments [38, 39, 40]. Here, ”environments” refer to distinct settings that can influence the structure and relationships within the graph data. These methods can be further classified into (1) invariant subgraph extraction, (2) adversarial learning, and (3) disentangled representation learning.
  • (1) Invariant Subgraph Extraction. Inspired by the invariant learning strategy in traditional OOD generalization tasks [41, 42, 43, 44], invariant subgraph extraction methods divide each graph into two segments [45, 46]: an invariant subgraph, which maintains stable structure-label relationships across distributions, and a variant subgraph, which may change across different distributions. By concentrating on the invariant subgraph, these methods leverage consistent relationships between graph structures and labels, which enhances the model’s ability to manage significant distribution shifts in OOD data, thereby improving its generalization performance. DIR [47] exemplifies this approach by distinguishing invariant subgraphs using a GNN-based approach, enabling the identification of consistent graph-label relations without the need for explicit distribution-based separation in the training data [47]. Conversely, GIL [45] advances this concept by utilizing the distributional divisions of training data to cluster variant subgraphs, thereby facilitating the generation of invariant graph representations for accurate classification. Additionally, MARIO [38] employs the graph Information Bottleneck (IB) principle [48, 49] for learning invariant representations by maximizing the mutual information between representations and labels while minimizing the mutual information between representations and inputs. Similarly, GSAT [46] formulates the graph IB by injecting learnable randomness into a node-level attention mechanism, which allows selecting subgraphs that keep invariant to the environmental shift and indicative to the labels.
    Recent work by Piao et al. [50] propose a hierarchical approach for effective subgraph extraction. This method generates multiple semantic environments for each graph while ensuring consistency across hierarchical levels. By considering multiple environments and modeling the correlations between different environments within a hierarchical framework, this approach can better understand the inherent variability in the data, leading to improved robustness against distribution shifts. Sun et al. [51] propose DIVE framework, which trains multiple models on all label-predictive subgraphs, consisting of both spurious and invariant subgraphs. This encourages the models to focus on distinct structural patterns within these subgraphs, facilitating the precise extraction of invariant subgraphs and mitigating the effects of simplicity bias.
  • (2) Adversarial Learning. Adversarial learning methods aim to learn invariant representations in different environments by enforcing models to perform consistently when adversarially enlarging the distribution shifts [52, 39]. Pioneering works of adversarial training adopt existing adversarial methods for OOD generalization from image data. For example, GNN-DRO [39] employs distributionally robust optimization [53] to address OOD challenges, aiming to minimize the worst-case loss within a defined range around the observed data distribution. However, the reliance of GNN-DRO on distribution labels can limit its application in the real world [54]. To address this limitation, GraphAT [40] introduces graph perturbations to create augmented distributions. These perturbations are achieved by maximizing the prediction discrepancies between a given node and its neighboring nodes, thereby effectively altering the graph’s smoothness to generate more diverse distributions.
  • (3) Disentangled Representation Learning. Disentangled representation learning aims to separate the distinct and informative factors in representations. In this manner, the learned disentangled representations contain decisive information for the target task. As a classic example, DisenGCN [55] introduces a multichannel convolutional layer along with a neighborhood routing strategy to achieve node representation disentanglement, and the disentangled representations are more robust to distribution shifts regarding node labels. IPGDN [56] enhances this disentanglement by promoting independence among latent factors based on Hilbert-Schmidt Independence Criterion (HSIC) [57], further improving generalization. From the perspective of factorization, FactorGCN [58] proposes to produce disentangled representations on the graph level by breaking down graphs into multiple interpretable factor graphs. As such, the graph-level generalization is strengthened. In addition to the graph-level generalization, NED-VAE [59] captures independent latent factors from both the node level and edge level with a variational autoencoder [60, 61]. This design improves generalization from various granularities of graph structures. Additionally, some works [62, 63, 64] point out that the degradation in OOD generalization ability often stems from spurious correlations between irrelevant patterns and category labels within the training data. To mitigate this, L2R-GNN [62] and OOD-GNN [63] utilize nonlinear graph decorrelation methods to eliminate spurious correlations between invariant patterns and other variables. L2R-GNN focuses on removing spurious correlations through a nonlinear approach, while OOD-GNN implements a global weight estimator for this purpose. Furthermore, EQuAD [65] proposes a graph invariance learning paradigm that disentangles invariant patterns from spurious features using the infomax principle [66, 67], thus enhancing the interpretability and robustness of learned representations.
    Another line of work has focused on utilizing contrastive learning to achieve disentanglement. For example, DGCL [68] employs a self-supervised strategy to learn disentangled representations through factor-wise contrastive learning, circumventing the high computational cost associated with self-supervised graph reconstruction. IDGCL [69] improves DGCL by introducing independence regularization to remove dependencies among disentangled representations. Different from other methods, IDGCL learns multiple disentangled representations, which are more interpretable and robust to distribution shifts.
    Beyond OOD generalization in conventional static graphs, dynamic graphs also face challenges from distribution shifts. Effectively modeling the complex latent environments and uncovering invariant patterns in dynamic graphs with distribution shifts remains an under-explored area. Zhang et al. propose DIDA [70], a framework that utilizes a disentangled spatio-temporal attention network to differentiate between variant and invariant patterns, thereby improving predictive stability in dynamic graphs under distribution shifts. I-DIDA [71] further refines predictive stability by identifying invariant patterns and applying disentangled spatio-temporal interventions. Similarly, Yuan et al. develop the EAGLE framework [72], which incorporates an environment-aware EA-DGNN to learn disentangled representations across various spatio-temporal environments. Zhou et al. [73] identify two key challenges in OOD generalization on spatiotemporal data: the entanglement of segment-level heterogeneity in temporal environments, and the failure of Invariant Risk Minimization (IRM) [42] on heterogeneous spatiotemporal observations. Their proposed CauSTG framework extracts invariant patterns through a spatiotemporal consistency learner and a hierarchical invariance explorer, directly addressing these challenges. Additionally, Wang et al. propose a STONE framework that employs a semantic graph learning module and a graph intervention mechanism capture invariant spatio-temporal dependencies among and thus address the spatio-temporal shifts [74].
Causality-based Learning. Rather than focusing on learning invariant or disentangled representations, causality-based learning methods focus more on uncovering the inherent causal information in graph data crucial for classification and mitigating spurious correlations that could potentially mislead classification. For instance, CIGA [11] explores the causal information within each graph by extracting essential causal components related to labels, thereby enhancing the generalizability of the model. DisC [33] adopts the GCE loss [75] to learn spurious information with a larger step size in order to capture the remaining causal information. To address the limitations of CIGA and DisC in scenarios where prior knowledge about the variance of causal and spurious information is unavailable, GALA [76] leverages an assistant model to detect distribution shifts in terms of such variance. It maximally learns the causal information based on proxy predictions, which helps in achieving successful OOD generalization under minimal assumptions. StableGNN [64] tackles spurious correlations in subgraph units through causal analysis. It incorporates a differentiable graph pooling layer to obtain subgraph-based representations and applies a causal variable distinguishing regularizer to eliminate spurious correlations. Recently, LECI [34] proposes a two-stage approach for environment-based methods, consisting of two stages: environmental inference that focuses on inferring environmental labels, and environmental exploitation that focuses on leveraging these labels for effective data utilization. In particular, LECI integrates causal independence principles into a subgraph discovery framework. Apart from the general OOD generalization problem caused by domain shifts, E-invariant GR [77] specifically investigates the issue of size shifts. It proposes a causal model that learns representations that are more robust against variations in graph sizes.
Graph Self-supervised Learning. Graph self-supervised learning has gained significant attention for its role in graph OOD generalization by facilitating the learning of more generalizable representations. Specifically, graph self-supervised learning enhances generalizability by mapping similar graphs to close representations [78]. In this manner, graph self-supervised learning improves the GNN models’ understanding of new test graphs, as their representations can closely align with the training data [79]. Pretraining-GNN [80] is among the first to explore various graph self-supervised learning methods at both the node and graph levels to pre-train GNN models, including context prediction, attribute masking, and structural similarity prediction. To address graph size shifts, PATTERN [27] introduces a self-supervised task focused on learning d-pattern representations from local structures, thereby enhancing generalization from smaller to larger graphs. Moreover, OOD-GCL [81] resorts to graph contrastive learning on factorized graph representations, thereby excluding the information from the latent factors sensitive to distribution shifts and improving overall generalization.
More recently, various Graph Foundation Models (GFMs) [82] have also been proposed based on graph self-supervised learning. GFMs exhibit superior generalizability on a variety of downstream tasks. One of the pioneering works in this area, GPPT [83] generalizes GNN models with prompt-tuning. Specifically, based on pre-trained GNN models via self-supervised learning, GPPT introduces additional tokens that represent candidate labels, thereby improving generalization to new tasks without the tedious fine-tuning of GNNs. Based on the prompt-tuning strategy, GPF [84] proposes a universal prompt-tuning framework to pre-train GNN models across various self-supervised learning strategies. Operating on the input graph’s feature space, GPF improves the generalizability of the prompted graph to various downstream tasks in an adaptive way. Drawing inspiration from ControlNet [85], GraphControl [86] enhances generalization with a universal structural pre-training strategy that enables the alignment of the input space across distributions and incorporates characteristics of target distributions as conditional contexts.
TABLE I: A summary of graph OOD generalization methods. ‘Task Level’ denotes the main task level, ‘Domain Source’ denotes whether the source domain consists of single or multiple domains, and ‘Distribution Shift’ denotes the distribution shifts that the methods aim to handle.
         Scenario          Approach          Technique          Method                    Task          Level                    Domain          Source                    Distribution          Shift
                   Graph OOD Generalization                    Model-centric                    Invariant          Structure Extraction          DIR [47]          Graph          Multiple          Covariate
         GIL [45]          Graph          Multiple          Covariate
         GSAT [46]          Graph          Multiple          Covariate
         IS-GIB [87]          Node/Graph          Single          Covariate
         MARIO [38]          Node          Multiple          Covariate&Concept
         DIVE [51]          Graph          Multiple          Covariate
                   Adversarial Learning          CAP [88]          Node          Single          Covariate
         GraphAT [40]          Node          Single          Covariate
         GNN-DRO [39]          Node          Single          Covariate&Concept
         WT-AWP [89]          Node/Graph          Single          Covariate
                   Disentangled          Representation Learning          DisenGCN [55]          Node          Single          Covariate
         IPGDN [56]          Node          Single          Covariate
         FactorGCN [58]          Node/Graph          Single          Covariate
         NED-VAE [59]          Node/Edge          Single          Covariate
         DGCL [68]          Graph          Single          Covariate
         IDGCL [69]          Graph          Single          Covariate&Concept
         I-DIDA [71]          Node/Edge          Single          Covariate&Concept
         L2R-GNN [62]          Graph          Single          Covariate
         OOD-GNN [63]          Graph          Single          Covariate
         EQuAD [65]          Graph          Multiple          Covariate
         DIDA [70]          Node          Single          Covariate
         EAGLE [72]          Node          Single          Covariate
         CauSTG [73]          Graph          Multiple          Covariate
         STONE [74]          Graph          Multiple          Covariate
                   Causality-based          Learning          E-invariant GR [77]          Graph          Single          Covariate (Size)
         CIGA [11]          Graph          Single          Covariate&Concept
         CAL [90]          Graph          Single          Covariate
         DisC [33]          Graph          Single          Covariate&Concept
         GALA [76]          Graph          Single          Covariate&Concept
         LECI [34]          Graph          Single          Covariate
         StableGNN [64]          Graph          Single          Covariate
                   Graph Self-supervised          Learning          Pretraining-GNN [80]          Node/Graph          Single          Covariate
         PATTERN [27]          Node          Single          Covariate
         OOD-GCL [81]          Node/Graph          Single          Covariate
         GPPT [83]          Node/Graph          Single          Covariate&Concept
         GPF [84]          Node/Graph          Single          Covariate&Concept
         GraphControl [86]          Node/Graph          Single          Covariate
                   Data-centric                    Graph Perturbation          AIA [91]          Graph          Single          Covariate
         G-Splice [35]          Node/Graph          Single          Covariate&Concept
         LiSA [37]          Node/Graph          Single          Covariate&Concept
         DLG [92]          Graph          Single          Covariate&Concept
         Pattern-PT [27]          Graph          Single          Covariate (Size)
         P-gMPNN [93]          Edge          Single          Covariate (Size)
                   Graph Mix-up          GraphMix [94]          Node          Single          Covariate
         G-Mixup [95]          Node/Graph          Single          Covariate
         $\mathcal{G}$-Mixup [96]          Graph          Single          Covariate
         OOD-GMixup [97]          Node/Graph          Multiple          Covariate
                   Distribution Augmentation          GREA[98]          Graph          Single          Covariate
         EERM [54]          Node          Multiple          Covariate
         FLOOD [99]          Node          Multiple          Covariate
         DPS [100]          Graph          Single          Covariate
         MoleOOD [101]          Graph          Single          Covariate
         ERASE [102]          Node          Single          Covariate
         IGM [103]          Graph          Multiple          Covariate
3.2 Data-Centric Approaches
It is widely accepted that the generalization ability of a model heavily relies on the diversity and quantity of the training data [104, 105]. Therefore, one intuitive yet efficient way to boost the model’s generalizability is to increase the diversity of training data through augmentation, including graph data augmentation and distribution augmentation.
Graph Data Augmentation. Graph data augmentation directly modifies graph structures and features to increase the diversity of existing training data, thereby exposing the model to a wider range of distributions to enhance its generalizability. Based on the used augmentation strategy, graph augmentation methods can be further classified into (1) graph prturbation and (2) graph mix-up.
  • (1) Graph Perturbation. For graph perturbation methods, original graphs are augmented by modifying the original graph structure such as adding or removing nodes and edges, as well as altering the representations of nodes and edges. In this manner, the augmented graphs are slightly different from the original graphs, thereby enhancing the diversity and quality of the training data. As a result, the model trained on the augmented data becomes more robust to various graph data distributions. Specifically, the goal of graph data augmentation methods for OOD generalization could be formulated as follows:
    $\min_{f,g}\mathbb{E}_{\textbf{A}^{\prime},\textbf{X}^{\prime},\textbf{Y}^{\prime}}[l(f(g(\textbf{A}^{\prime},\textbf{X}^{\prime})),\textbf{Y}^{\prime})],$ (4)
    where $\textbf{A}^{\prime}$, $\textbf{X}^{\prime}$, and $\textbf{Y}^{\prime}$ represent the adjacency matrix, node feature matrix, and label of a perturbed graph, respectively. In this formulation, $g$ transforms the perturbed graph inputs into latent representations, while $f$ functions as the classifier.
    As a classic example, Pattern-PT [27] studies the problem of graph size generalization and discovers that the discrepancy in small local structures across graphs could harm the generalizability of GNNs. To mitigate this issue, Pattern-PT proposes an interpolation strategy among graphs in order to learn comprehensive representations of local structures. P-gMPNN [93] focuses on the link-level OOD generalization problem with graph size shifts. This work explores the vulnerability of GNNs as link predictors based on structural node embeddings, and proposes a theoretically sound method that learns robust pairwise embeddings by modifying graphs into graphon fractions of common neighbors. To avoid causing significant changes and losing relevant information during augmentation, several works aim to enhance the stability of augmentation strategies. For instance, AIA [91] considers the causal invariance principle while enlarging the diversity of environments, aiming to reduce the distribution gap between the augmented environments and the potential test data unseen during training. To ensure causal invariance, AIA designs a causal generator and adversarially learns the masks that encapsulate causal information. Similarly, G-Splice [35] is proposed to avoid the corruption of the underlying causal information related to labels. The authors explore the extrapolation strategies from both the perspectives of structures and features to augment graph data without impacting the causal information. From the perspective of label shifts, LiSA [37] identifies that augmented graph labels may change during structural modifications, and hence proposes to avoid such a label shift issue during graph augmentation. Particularly, LiSA employs a variational subgraph generator, trained to produce label-invariant subgraphs, which are subsequently used in aiding graph OOD generalization. DLG [92] improves the label and distribution consistency by maximizing the mutual information between augmented and original graphs.
  • (2) Graph Mix-up. The mixup strategies, initially proposed and shown to be effective in computer vision and natural language processing, have demonstrated significant advantages in enhancing generalization through interpolating instances [106, 107, 108]. These techniques work by creating synthetic training graph data through mix-up, which enhances model robustness to distribution shifts and shows promise in graph OOD generalization [109]. For instance, GraphMix [94] integrates manifold mixup [110] for generalizable node classification by simultaneously training a fully-connected network (FCN) and a GNN, with a parameter-sharing strategy that facilitates the transfer of essential node representations from the FCN to the GNN. Similarly, G-Mixup [95] extends this concept by interpolating latent node representations in the embedding space, offering generalizability across different graphs. Recognizing that latent representations may lack richness, NodeAug [111] and ifMixup [112] further propose to directly interpolate raw node features and graph data, respectively, to create more diverse inputs for GNNs.
    On the other hand, $\mathcal{G}$-Mixup [96] proposes to perform mixup on graphons across classes in the Euclidean space, resulting in new graphs derived from the interpolation of graphons. The results demonstrate enhanced performance in graph classification tasks with distribution shifts and highlight the efficacy $\mathcal{G}$-mixup in promoting OOD generalization. Instead of relying on predefined disturbances, OOD-GMixup [97] generates virtual samples by applying manifold mixup directly to the graph rationale representations. This approach allows for the simulation of complex graph distribution shifts by manipulating these representations within a metric space. Furthermore, to accurately measure and control the training process amid these shifts, OOD-GMixup introduces an OOD confidence score, derived using Extreme Value Theory (EVT). This score, coupled with a sample re-weighting mechanism, actively manages the training dynamics by prioritizing virtual OOD samples, thereby enhancing the model’s generalization capabilities across varied and unforeseen graph distributions.
    While extensive research efforts [96, 94, 95] have focused on interpolation for synthetic data augmentation, relatively few studies have explored the extrapolation in addressing out-of-distribution challenges on graphs. Recently, a novel framework [113] was proposed to generate synthetic graphs using a non-Euclidean-space linear extrapolation method. The generated graphs have been theoretically validated to adhere to common causal assumptions. The framework consists of two components, which are designed to generate synthetic graphs in both the structural and feature spaces. Extensive experiments demonstrate that this extrapolation approach enhances a model’s OOD generalization ability on graphs and outperforms previous invariant learning methods [45, 47, 11] and other graph data augmentation techniques [96, 94, 95].
Distribution Augmentation. In contrast to graph data augmentation, which enhances the original graph at the graph level, distribution augmentation focuses on generating new distributions of graphs to facilitate robust representation learning [54]. Distribution augmentation is particularly useful when the distribution available during training is scarce, making learning robust representations extremely challenging, as the model struggles to distinguish between decisive information and spurious information that varies across distributions.
GREA [98] is a pioneering augmentation approach to compensate for the lack of sufficient distribution. It operates under the assumption of the graph rationale, i.e., for each input graph, there exists a specific subgraph that is decisive for its prediction. Therefore, GREA proposes to augment each graph by replacing the variant subgraph (i.e., subgraphs that do not involve the graph rationale) of the target graph with the variant subgraphs from other existing graphs. In this way, by training on these augmented graphs with varying domain subgraphs, the models will learn to focus solely on the graph rationale for making inferences, thereby enhancing OOD generalization. On the other hand, DPS [100] develops and utilizes several subgraph generators to generate diverse augmented subgraphs as additional distribution, while preserving the subgraphs with graph rationale. An additional regularization term is further introduced to amplify the difference among augmented distributions, enhancing the diversity of distributions of model training in order to improve the generalizability of the model. Following this, IGM [103] identifies invariant subgraphs and generates diverse environments by concatenating these with environment-related subgraphs. This targeted mixing approach not only helps eliminate spurious correlations but also fosters sufficient distribution shifts essential for robust invariant learning.
Additionally, EERM [54] proposes to generate new distribution by leveraging invariance principles for prediction. EERM employs multiple context explorers, namely the graph structure editors, which are adversarially trained to maximize the variance of risks across multiple generated distributions. Such a design helps address the challenge of generating new distribution caused by the inter-connections among nodes on graphs. Based on EERM, FLOOD [99] further incorporates a bootstrapped learning module that could be refined on the test data to improve OOD generalization. ERASE [102] enhances generalization by employing a decoupled label propagation approach that corrects noisy labels through structural denoising before training. This method ensures that the model learns from more accurate representations of the data, effectively distinguishing meaningful information from noise. By maximizing coding rate reduction during training, ERASE adapts to diverse data distributions, improving robustness against label noise.
4 Training-Time Graph OOD Adaptation
Training-time graph OOD adaptation focuses on addressing the distribution discrepancies between source and target graph data. By mitigating the distribution shifts, models can effectively transfer knowledge from the source domain and adapt to the target distribution. In general, training-time graph OOD adaptation serves three primary objectives in different scenarios.
  • Observation Bias Correction. For semi-supervised classification within a single graph, distribution shifts between training and test instances may arise from observation bias related to latent subpopulation [4], or the time-evolving nature of graphs [114]. Mitigating distribution shifts in this setting may enhance the model’s performance on test instances.
  • Cross-graph Knowledge Transfer. To facilitate knowledge transfer from well-labeled graphs to those with limited labels, it is essential to effectively manage distribution shifts between these graphs, as distinct graphs often exhibit varied data distributions.
  • Negative Augmentation Mitigation. Graph data augmentation, which utilizes the augmented data as additional training data, is commonly used for improving model generalization or alleviating label scarcity issues. However, overly severe distribution shifts between the original and augmented data can lead to the negative augmentation problem [115, 116]. Therefore, controlling distribution shifts is essential to avoid inferior model performance and fully exploit the benefits of augmented graph data.
In this section, we discuss existing training-time graph OOD adaptation methods, highlighting the techniques for mitigating distribution shifts. The summary can be found in Table II.
4.1 Model-Centric Approaches
Model-centric methods for training-time graph OOD adaptation can be further categorized into invariant representation learning and concept-shift aware representation learning, which aim at learning distributionally aligned representations, and model regularization, which focuses on achieving effective knowledge transfer through model regularization.
Invariant Representation Learning. Similar to the invariant representation learning adopted in graph OOD generalization, the goal of this technique in training-time graph OOD adaptation is to learn invariant representations across source and target graphs, ensuring robustness to distribution shifts. This technique is frequently employed for domain adaptation under covariate assumption [^1] Within the context of graph OOD adaptation, we commonly use the terms “distribution” and “domain” interchangeably., in which an invariant relationship between latent representations and labels $\mathcal{P}_{S}(\textbf{Y}|\textbf{H})=\mathcal{P}_{T}(\textbf{Y}|\textbf{H})$ is assumed. Inspired by the theoretical generalization bound [117], domain-invariant representation learning methods aim to train a graph encoder $g(\cdot)$ to minimize discrepancies between the induced marginal source distribution $\mathcal{P}_{S}(\textbf{H})$ and the target distribution $\mathcal{P}_{T}(\textbf{H})$, while also identifying a classifier $f(\cdot)$ in the latent space that minimizes empirical source risk. To achieve these two goals, the loss function for domain-invariant representation learning can be formulated as follows:
$\min_{f,g}\mathbb{E}_{\textbf{A, X, Y}}[l(f(g(\textbf{A},\textbf{X})),\textbf{Y})]+l_{reg},$ (5)
where $l_{reg}$ denotes a regularization term that facilitates the alignment of the induced marginal distribution $\mathcal{P}(\textbf{H})$. Three strategies are mainly adopted: (1) distribution distance minimization, (2) adversarial learning, and (3) disentangled representation learning.
  • (1) Distribution Distance Minimization. Distribution distance minimization directly employs the distance between marginal distributions as the regularization term in Equation 5. Methods vary in terms of the choice of the distance metric and the specific representations they aim to align. SR-GNN [118] considers central moment discrepancy as regularization and aligns the distribution discrepancies in the final layer of traditional GCN. CDNE [119], GraphAE [12], and GRADE [120] target at minimizing the statistical discrepancies between source and target across all latent layers, with the regularization term as a summation of distribution distances of different layers. In terms of their distance metric, CDNE uses the marginal maximum mean discrepancy and class-conditional marginal maximum mean discrepancy, GraphAE considers the multiple kernel variant of the maximum mean discrepancy as the distance metric, and GRADE defines and utilizes subtree discrepancy. JHGDA [121] relies on a hierarchical pooling module to extract network hierarchies and minimizes statistical discrepancies in hierarchical representations via the exponential form of the marginal and class-conditional maximum mean discrepancy. For non-trainable representations, such as the latent embeddings in SimpleGCN [2], SR-GNN [118] applies an instance weighting technique. In this approach, the learnable weight parameters are optimized using kernel mean matching to mitigate distribution discrepancies. HC-GST [122] uses pseudo-labels, incorporating heterophily ratios and labeling accuracy to align distributions, with learnable weights optimized via kernel mean matching. Most recently, DREAM [123] employs complementary branches (graph-level and subgraph-enhanced) and enforces consistency between them to mitigate biased pseudo-labels and reduce distribution discrepancies. Addtionally, SelMAG [124] designs an optimal-transport-based algorithm by selecting informative source graphs and adapting both embedding and classification spaces, leveraging meta-learning for precise sub-graph and node selection to improve transferability.
  • (2) Adversarial Learning. Adversarial learning approaches align representations by training the graph encoder $g(\cdot)$ to generate embeddings that confuse the domain discriminator $f_{d}(\cdot)$. Correspondingly, the regularization term in Equation 5 is usually formulated as a minimax game between $g(\cdot)$ and $f_{d}(\cdot)$ as:
    $\min_{f_{d}}\max_{g}l(f_{d}(g(\textbf{A},\textbf{X})),\textbf{Y}_{d}),$
    where $Y_{d}$ denotes the domain label, and $l$ can be chosen as a negative distance loss [125], or a domain classification loss [126, 127, 128, 129, 12, 13, 29]. In addition to framing it as a minimax problem, DANE [130] explores the use of two symmetric and adversarial losses for model training, aiming to achieve bi-directional transfer. Typically, adversarial alignment takes place in the final hidden layer, with the exception being GraphAE [12], which aligns representations in all hidden layers. In addition, it is noteworthy that SGDA [13] also takes the label scarcity issue of the source graph into account by employing a weighted self-supervised pseudo-labeling loss. SDA [29] studies the problem under the open-set setting, utilizing an entropy-based separation strategy to categorize target nodes into certain and uncertain groups and specifically aligning nodes from the certain group with adversarial learning technique. Most recently, JDA-GCN [131] introduces a joint adversarial domain adaptive graph convolutional network, leveraging both local and global graph structures through structural graph alignment, improving the model’s ability to capture intricate dependencies within graph data.
  • (3) Disentangled Representation Learning. Disentangled representation learning separates representations into distinct, informative factors. Among these factors, one is designed to capture domain-invariant features that are essential for the target task (e.g classification). This domain-invariant component remains consistent across different domains to preserve classification-related semantic information, thereby facilitating effective adaptation. And the disentangled components can be selectively manipulated to cater to specific requirements of the target task. The loss function for disentangled representation learning can be defined as:
    $\min_{f,g_{s}}\mathbb{E}_{\textbf{A, X, Y}}[l(f(g_{s}(\textbf{A},\textbf{X})),\textbf{Y})]+\min_{g_{s},g_{o}}(l_{reg}+l_{recon}+l_{add}),$
    where $g_{s}$ denotes the graph encoder for acquiring invariant task-related information, $g_{o}$ denotes graph encoder(s) for other components excluding $g_{s}$, $l_{reg}$ represents a regularization term for enhancing the separation between different components, and $l_{recon}$ denotes a reconstruction loss aiming to recover the original graph structure from the concatenated representation, thereby preventing information loss. Additional terms $l_{add}$ are introduced to facilitate the learning of disentangled representations, enabling specific components to exhibit the desired characteristics. In ASN [132], the representation is decomposed into a domain-private part and a domain-invariant classification-related part. A domain adversarial loss is additionally added to facilitate the learning of invariant representations. Analogous to DIVA [133], DGDA [134] assumes that the graph generation process is controlled independently by domain-invariant semantic latent variables, domain latent variables, and random latent variables. To learn representations with desired characteristics, domain classification loss and noise reconstruction loss are considered as the additional losses.
Concept-shift Aware Representation Learning. Some recent works extend beyond the covariate assumption and take the change of label function across domains into consideration. As illustrated in [135], when there exist concept shifts in $\mathcal{P}(\textbf{Y}|\textbf{X})$ or $\mathcal{P}(\textbf{Y}|\textbf{H})$, namely, the label function changes, the inestimable adaptability term in the upper bound [117] may be large and the performance of previous invariant representation learning methods on target is no longer guaranteed. A similar upper bound and example provided in [5] illustrate the insufficiency of invariant representation learning methods based on the covariate shift assumption.
To further accommodate the change in label function, SRNC [114] leverages graph homophily, incorporating a shift-robust classification GNN module and an unsupervised clustering GNN module to alleviate the distribution shifts in the joint distribution $\mathcal{P}(\textbf{H},\textbf{Y})$. Notably, SRNC is also capable of handling the open-set setting where new classes emerge in the test data. In StruRW [5], the conditional structure shifts $\mathcal{P}(\textbf{A}|\textbf{Y})$ are identified and mitigated by adaptively adjusting the weights of edges in the source graph. Building on this, Pair-Align [136] is proposed as an extension of StruRW to address both conditional structure shift and label shift. Moreover, Zhu et al. [17] demonstrate that, in the contextual stochastic block model, conditional shifts in the latent space $\mathcal{P}(\textbf{Y}|\textbf{H})$ can be exacerbated by both graph heterophily and graph convolution in the GCN compared to conditional shifts in input feature space $\mathcal{P}(\textbf{Y}|\textbf{X})$. Hence, they introduce GCONDA that explicitly matches the distribution of $\mathcal{P}(\textbf{Y}|\textbf{H})$ across domains via Wasserstein distance regularization, and additionally, they also propose GCONDA++ that jointly minimizes the discrepancy in $\mathcal{P}(\textbf{Y}|\textbf{H})$ and $\mathcal{P}(\textbf{H})$.
Model Regularization. Instead of focusing on the process of learning aligned representations, some other methods achieve effective knowledge transfer under distribution shifts through model regularization strategy, including both spectral regularization [16, 137] and knowledge distillation [115, 4]. Building on the derived GNN-based generalization bound, You et al. [16] propose SSReg and MFRReg, which regularize the spectral properties of GNN to enhance transferability. They also extend their theoretical results to the semi-supervised setting with the challenging distribution shifts in $\mathcal{P}(\textbf{Y}|\textbf{A},\textbf{X})$. Liu et al. [137] then delve into the inherent generalization ability of GNN models, proposing the removal of propagation layers in the source graph while stacking multiple propagation layers in the target graph. This model design strategy inherently acts as a regularization for the GNN Lipschitz constant and helps to tighten the generalization bound. Both KDGA [115] and KTGNN [4] employ knowledge distillation strategy and regularize the Kullback–Leibler (KL) divergence between the outputs of teacher and student models. In particular, KDGA aims to address the negative augmentation problem by distilling the knowledge of a teacher model trained on augmented graphs into a partially parameter-shared student model on the original graph. KTGNN, on the other hand, considers the semi-supervised node classification problem for VS-Graph, in which vocal nodes are regarded as the source and silent nodes with incomplete features are regarded as the target. The source classifier and target classifier are respectively constructed, and the knowledge of both of the classifiers is distilled into the student transferable classifier through the KL regularization.
TABLE II: A summary of training-time and test-time graph OOD adaptation methods.
Scenario Approach Technique Method Task Level Domain Source Distribution Shift Objective
Training-time Graph OOD Adaptation Model-centric Invariant Representation Learning DAGNN [126] Graph Single Covariate Cross-graph transfer
DANE [130] Node Single Covariate Cross-graph transfer
CDNE [119] Node Single Covariate&Concept Cross-graph transfer
ACDNE [128] Node Single Covariate Cross-graph transfer
UDA-GCN [127] Node Single Covariate Cross-graph transfer
DGDA [134] Graph Single Covariate Cross-graph transfer
SR-GNN [118] Node Single Covariate Observation bias correction
COMMANDER [129] Graph Single Covariate Cross-graph transfer
ASN [132] Node Single Covariate Cross-graph transfer
AdaGCN [125] Node Single Covariate&Concept Cross-graph transfer
GraphAE [12] Node Single/Multiple Covariate Cross-graph transfer
GRADE [120] Node/Edge Single Covariate Cross-graph transfer
JHGDA [121] Node Single Covariate Cross-graph transfer
SGDA [13] Node Single Covariate Cross-graph transfer
SDA [29] Node Single Covariate&Concept Cross-graph transfer
JDA-GCN [131] Node Single Covariate Cross-graph transfer
HC-GST [122] Node Single Covariate Observation bias correction
DREAM [123] Graph Single Concept Cross-graph transfer
SelMAG [124] Node Multiple Covariate&Concept Cross-graph transfer
Concept-shift Aware Representation Learning SRNC [114] Node Single Covariate&Concept Cross-graph transfer/Observation bias correction
StruRW [5] Node Single/Multi Concept (Structure) Cross-graph transfer
Pair-Align [136] Node Single Covariate&Concept Cross-graph transfer
GCONDA++ [17] Node/Graph Single Covariate&Concept Cross-graph transfer/Observation bias correction
Model Regularization KDGA [115] Node Multiple Concept (Structure) Negative augmentation mitigation
SS/MFR-Reg [16] Node/Edge Single Covariate&Concept Cross-graph transfer
A2GNN [137] Node Single Covariate Cross-graph transfer
KTGNN [4] Node Single Concept (Feature) Observation bias correction
Data-centric Instance weighting IW [138] edge Single Concept (Struture) Cross-graph transfer
NES-TL [139] Node Multiple Concept (Struture) Cross-graph transfer
RSS-GCN [140] Graph Multiple Covariate Cross-graph transfer
DR-GST [116] Node Single Covariate&Concept Negative augmentation mitigation
Graph Data Augmentation FakeEdge [141] Edge Single Concept (Structure) Observation bias correction
Bridged-GNN [142] Node Single Covariate&Concept Cross-graph transfer
DC-GST [143] Node Multiple Covariate Negative augmentation mitigation/Bias correction
LTLP [144] Edge Single Covariate Negative augmentation mitigation/Cross-graph transfer
Test-time Graph OOD Adaptation Model-centric Semi-supervised Fine-tuning GraphControl [19] Node Multiple Concept (Feature) Avoid over-fitting
G-Adapter [145] Graph Single Covariate&Concept Avoid over-fitting
AdapterGNN [146] Graph Single Covariate&Concept Avoid over-fitting
PROGRAM [147] Node Multiple Covariate&Concept Avoid over-fitting
Self-supervised Adaptation SOGA [148] Node Single Covariate Maintain discriminative ability
GAPGC [149] Graph Single Covariate Maintain discriminative ability
GT3 [14] Graph Multiple Covariate Maintain discriminative ability
RNA [150] graph Single Concept Maintain discriminative ability
Data- centric Feature Reconstruction FRGNN [151] Node Single Concept (Feature) Mitigate bias
Graph Data Augmentation GTRANS [18] Node Single Covariate Maintain discriminative ability
GraphCTA [15] Node Single Covariate Maintain discriminative ability
SGOOD [152] Graph Multiple Covariate Maintain discriminative ability
4.2 Data-Centric Approaches
In addition to model-centric strategies, data-centric approaches play a crucial role in addressing training-time graph OOD adaptation challenges. These methods focus on leveraging and enhancing graph data itself to improve adaptation performance. By considering the unique characteristics of graphs, these approaches can effectively address the complexities introduced by distribution shifts. Two prominent data-centric techniques are instance weighting and graph data augmentation.
Instance Weighting. Instance Weighting, which assigns different weights for data points, is a commonly used data-centric technique in traditional transfer learning [8]. Similar strategies of weighting graph instances are also adopted in methods for training-time graph OOD adaptation, including edge weighting [138], node weighting [116], and graph weighting [139, 140] for specific tasks or purposes. Borrowing the idea from Adaboost and TrAda, Ye et al. [138] employ the instance weighting technique for the edge sign prediction task. Edge weights are adjusted in each iteration, with reduced weights assigned to misclassified dissimilar source instances to mitigate distribution shifts across graphs. On the other hand, Liu et al. [116] recognize that the distribution shifts between the original data and the augmented data with pseudo-labels may impede the effectiveness of self-training, and thus propose to reweight the augmented node instances based on information gain to mitigate the gap between the original distribution and the shifted distribution. In multi-source transfer settings, the source graphs may not be equally important for predictions on the target graph and some of them may be of poor quality. To address such challenges, both RSS-GCN [140] and NES-TL [139] employ graph weighting techniques to effectively combine the available source graphs. NES-TL [139] proposes the NES index to quantitatively measure the structural similarity between two graphs, and use the NES-based scores as weights to ensemble weak classifiers trained on instances from each source graph and labeled target instances. On the other hand, RSS-GCN [140] utilizes reinforcement learning to select high-quality source graphs, aiming to minimize the distribution divergence between selected source graphs and target graphs. Such sample selection strategy can be considered as a special binary instance weighting.
Graph Data Augmentation. Different from the aforementioned weighting strategy, several works have delved into exploring graph data augmentation strategies to mitigate distribution shifts, mainly focusing on edge adaptation [141, 142, 143]. Dong et al. [141] find that the dataset shift challenge in edge prediction arises from the presence of links in training and the absence of link observations in testing. To address this, they propose FakeEdge, a subgraph-based link prediction framework that intentionally adds or removes the focal links within the subgraph. This approach aims to decouple the dual role of edges, functioning both as components in representation learning and as labels for link prediction. Bi et al. [142] reconsider the domain-level knowledge transfer problem as learning sample-wise knowledge-enhanced posterior distribution. They first learn the similarities of samples from both source and target graphs and build bridges between each sample and its similar samples containing valuable knowledge for prediction. A GNN model is then employed to transfer knowledge across source and target samples on the constructed bridged-graph. More recently, a novel framework called DC-GST [143] has been introduced to bridge the distribution shifts between augmented training instances and test instances in self-training, which incorporates a distribution-shift-aware edge predictor to improve the model’s generalizability of assigning pseudo-labels, along with a carefully designed criterion for pseudo-label selection. Additionally, LTLP [144] enhances graph data augmentation by addressing the long-tailed distribution of common neighbors in link prediction, effectively increasing the common neighbors of tail node pairs through high-quality edge generation and representation refinement.
5 Test-Time Graph OOD Adaptation
Test-time graph OOD adaptation focuses on adjusting a pre-trained model to new target data that becomes available after the initial training phase. Unlike training-time adaptation, which requires both source and target data simultaneously, test-time adaptation adjusts the model using only the target data. This approach, often referred to as source-free adaptation, plays a crucial role in scenarios where access to source data is restricted. For example, in social networks, source data is typically confidential and inaccessible due to privacy protection purposes and data leakage concerns. In this section, we delve into test-time graph OOD adaptation. A summary for related models is provided in Table II.
5.1 Model-Centric Approaches
Fine-tuning is a widely used approach to adapt a pre-trained model during the test time. However, effectively leveraging information from the pre-trained model presents challenges, which are generally addressed by the following two techniques:
Semi-supervised Fine-tuning. In this technique, the model is usually pre-trained in an unsupervised way to encode more transferable and generalizable structural information. Subsequently, during fine-tuning, task-related and domain-specific information is incorporated. A notable challenge here is the scarcity of labels in the target graphs, which can lead to overfitting. To avoid negative transfer, both G-Adapter [145] and AdapterGNN [146] employ the parameter-efficient fine-tuning (PEFT) strategy that takes bottleneck structure to alleviate overfitting by reducing the size of the tunable parameter. Alternatively, GraphControl [19] considers adaptively integrating the target information, incorporating two components: a frozen pre-trained model that takes the adjacency matrix as input and a trainable copy that takes node feature-based kernel matrix as input. The two components are connected through zero MLPs with expanding parameters, aiming to mitigate noise in target node features while gradually integrating downstream information into the pre-trained model. PROGRAM [147] builds robust connections between test data and prototypes with two key components: a prototype graph model that generates reliable pseudo-labels by leveraging relationships between prototypes and test data, and a robust self-training module that refines these pseudo-labels iteratively through consistency regularization.
Self-supervised Adaptation. In this technique, task-related information is encoded into the pre-trained model, and subsequently, an unsupervised fine-tuning task on the target graph is followed. However, a major challenge arises during this unsupervised task: the model may lose its discriminative ability for the primary task or potentially learn irrelevant information. To address this issue, various techniques have been proposed to retain the model’s discriminative power. SOGA [148] and GAPGC [149] employ information-based design strategies while GT3 [14] tackles the challenge by intentional architecture design and additional regularization constraints. In SOGA, Mao et al. [148] utilized a loss that maximizes the mutual information between the inputs and outputs of the model to enhance the discriminatory power, whereas GAPGC [149] employs an adversarial pseudo-group contrast strategy to tackle over-confidence bias and mitigate the risk of capturing redundant information, which has a lower bound guarantee of the information relevant to the main task from the information bottleneck perspective. On the other hand, GT3 [14] structures the model to include two branches that share initial layers: a main task (classification) branch and a self-supervised branch, and only the self-supervised branch is tuned during test time so that the discriminative ability in the main-task branch can be maintained. Additionally, it also integrates additional regularization constraints between training and test output embeddings, hence enforcing their statistical similarity and avoiding substantial fluctuations.
Other relevant studies, such as GTOT-Tuning [153], which transfers knowledge across tasks during test time, and GraphGLOW [154], which incorporates sharable components directly into the model, also provide valuable insights into addressing distribution shifts during test-time adaptation.
5.2 Data-Centric Approaches
Data-centric approaches in test-time graph OOD adaptation focus on modifying the input data to effectively address distribution shifts between the training and target domains. These methodologies emphasize the reconstruction and augmentation of graph features to enhance the model’s performance without altering the pre-trained architecture.
Feature Reconstruction. Feature reconstruction aims to mitigate distribution shifts between training and target through reconstructing node features of target graphs to resemble those of source graphs during test time. In FRGNN, Ding et al. [151] consider the semi-supervised node classification task and utilize an MLP to establish a mapping between the output and input space of the pre-trained GNN. Subsequently, using the encoded one-hot class vectors as inputs, the MLP generates class representative representations. By substituting the features of the labeled test nodes with the representative representations of the corresponding classes and spreading the updated information to other unlabeled test nodes through message passing, the graph embedding bias between test nodes and training nodes is anticipated to be mitigated.
Graph Data Augmentation. Apart from reconstructing features, Jin et al. [18] introduce a graph transformation framework called GTRANS to address distribution shifts during test time. The graph transformation is modeled as injecting perturbations on the graph structure and node features, and is subsequently optimized via a parameter-free surrogate loss. Theoretical analyses are also provided to guide the selection of appropriate surrogate loss functions. Upon GTRANS, GraphCTA [15] further considers collaboratively integrating model adaptation and graph adaptation. The collaborative loop is formed by adapting the model using predictions from the node’s neighborhood, enhancing the graph through neighborhood contrastive learning, and then reintroducing the adapted graph into the model adaptation process. Additionally, the SGOOD framework [152] leverages task-agnostic substructures to enhance graph-level OOD detection by encoding their relationships into effective representations It is worth highlighting that such data-centric test-time graph OOD adaptation methods that prioritize adjusting the test data over modifying the pre-trained model can be especially beneficial when handling large-scale pre-trained models.
6 Dataset
TABLE III: A summary of graph OOD datasets. ”Avg. Nodes” and ”Avg. Edges” represent the average number of nodes and edges in each graph, respectively. ”Classes” indicates the number of target categories for classification tasks. ”Graphs” specifies the total number of graphs in each dataset. ”Scenario” refers to the primary context of generalization (G.) or adaptation (A.), while ”Shift Types” denote the types of distribution shifts, with covariate (Cov.) and concept (Con.) shifts. ”Domain Separation” identifies the factors that separate the domains. ”Field” describes the nature of the data, and ”Task” highlights the level at which the dataset can be used.
Dataset Avg. Nodes Avg. Edges Classes Graphs Scenarios Shift Types Domain Separation Field Task Level
Spurious-Motif [47] 44.96 65.67 3 18000 G. Cov.&Con. Size Synthetic graph Graph
MNIST-75sp [155] 70.97 590.52 10 35000 G. Cov.&Con. Feature noise Superpixel graph Graph
Graph-SST2 [156] 10.2 18.4 2 43689 G. Cov. Node degree Text sentiment Graph
DrugOOD-Assay [157] 32.27 70.25 2 72239 G. Cov. Assay Molecular graph Graph
DrugOOD-Scaffold [157] 29.95 64.86 2 59608 G. Cov. Scaffold Molecular graph Graph
DrugOOD-Size [157] 30.73 66.90 2 70672 G. Cov. Size Molecular graph Graph
D&D [158] 284.32 1431.32 2 1178 G. Cov. Graph size Molecular graph Graph
NCI1 [158] 29.87 32.30 2 4110 G. Cov. Graph size Molecular graph Graph
OGBG-MolPCBA [159] 26.0 28.1 2 437929 G. Cov.&Con. Scaffold/Size Molecular graph Graph
OGBG-MolHIV [159] 25.5 27.5 2 41127 G. Cov.&Con. Scaffold/Size Molecular graph Graph
IMDB [160] 19.77 96.53 2 1000 G. Cov. Time Collaboration network Graph
Twitter [161] 0.91 2.51 2 499160 A. Cov. Time Social network Graph
Reddit-Binary [3] 429.63 497.75 2 2000 G./A. Con. Community Social network Graph/Node
OGBG-PPA [159] 243.4 2266.1 37 158110 G. Cov.&Con. Species Protein network Graph
OGBN-Proteins [159] 132534 39561252 2 1 G. Cov. Species Protein network Node
PPI [3] 2372.7 68508.7 24 24 G. Cov.&Con. Graph Protein network Node/Edge
OGBN-Products [159] 2449029 61859140 47 1 G. Cov. Popularity Co-purchase network Node
Amazon-Photo [162] 7650 238162 8 1 G./A. Cov. Graph Co-purchase network Node
Cora [163] 2708 5429 7 1 G./A. Cov.&Con. Word/Degree Citation network Node
Citeseer [164] 3327 4732 6 1 G./A. Cov.&Con. Word/Degree Citation network Node
Pubmed [165] 19717 44338 3 1 G./A. Cov.&Con. Word/Degree Citation network Node
OGBN-Arxiv [159] 169343 1166243 40 1 G./A. Cov.&Con. Time Citation network Node
DBLP [17] 39254.5 500650 5 2 G./A. Cov.&Con. Time Citation network Node
ACM [17] 11671.5 81053 5 2 G./A. Cov.&Con. Time Citation network Node
Yelp [124] 2317.17 38754.17 6 6 G./A. Cov.&Con. City Social network Node/Edge
Facebook100 [166] 12066.08 469028.26 2 100 G. Cov.&Con. University Social network Node
Twitch [167] 5686.67 148724.33 2 6 G. Cov.&Con. Language Gamer network Node
Elliptic [168] 203769 234355 2 1 G./A. Con. Time Bitcoin Transactions Node
GOOD-WebKB [169] 617 1138 5 1 G. Cov.&Con. University Webpage network Node
BlogCatalog [170] 5196 171743 6 2 A. Cov.&Con. Group affiliation Blog network Node
USA-Airport [171] 1190 13599 4 1 A. Cov.&Con. Country Air-traffic network Node
Brazil-Airport [171] 131 1038 4 1 A. Cov.&Con. Country Air-traffic network Node
Europe-Airport [171] 399 5999 4 1 A. Cov.&Con. Country Air-traffic network Node
Yeast [158] 21.54 22.84 2 79601 A. Cov. Species Protein network Node/Edge
OGBL-PPA [159] 576289 30326273 - 1 G. Cov. Throughput Protein network Edge
OGBL-COLLAB [159] 235868 1285465 - 1 G. Con. Time Collaboration network Edge
Evaluating models under distribution shifts in graph learning is crucial for assessing their generalizability and adaptability. In this section, we introduce common datasets and benchmarks that are typically used for such tasks, covering a range of fields and types of distribution shifts.
Open Graph Benchmark (OGB) [159]. OGB is a widely used platform that provides a collection of standardized datasets for machine learning on graphs. It includes diverse task types such as node classification, graph classification, and link prediction, covering large-scale molecular graphs, biological networks, and social networks. OGB datasets are often used to study how models generalize under distribution changes or domain shift scenarios. Notable datasets such as OGBG-MolPCBA and OGBG-PPA provide molecular and protein interaction networks, respectively, and are commonly utilized for OOD generalization and adaptation tasks.
TUdataset [158]. TUdataset is another popular collection of benchmark datasets, including graph classification and regression tasks across different domains, such as bioinformatics (NCI1, D&D) and social networks (IMDB, Reddit). These datasets are frequently used to evaluate model generalization across graphs of varying sizes and types, making them useful for cross-domain tasks that involve distribution shifts in graph structures and attributes.
GOOD Benchmark (Graph Out-of-Distribution Benchmark) [169]. The GOOD benchmark is specifically designed to assess the performance of graph learning models under out-of-distribution (OOD) scenarios. GOOD includes several tasks where models need to handle shifts in graph data distribution, such as covariate and concept shifts. For example, the Spurious-Motif dataset in GOOD challenges models to discern between spurious and true features within graphs, offering a robust test for generalization under distribution shifts.
DrugOOD [157]. DrugOOD focuses on OOD challenges in AI-aided drug discovery, particularly affinity prediction tasks with noisy annotations. This benchmark includes molecular graph datasets and evaluates model performance across various settings, such as Assay, Scaffold, and Size splits, where domain-specific knowledge is crucial for generalization under distribution shifts. DrugOOD primarily considers covariate shifts, thus providing insights into domain generalization for molecular graphs.
Notably, most existing graph learning benchmarks include only one type of shift. In addition to the benchmarks mentioned above, other widely-used datasets in graph learning tasks under distribution shifts include citation networks like Cora, Citeseer, and Pubmed, which are commonly used for node classification tasks. Table III summarizes the commonly used datasets for graph learning under distribution shifts, detailing their properties such as the average number of nodes, edges, classes, and graphs, as well as the applicable task types and distribution shift scenarios. These datasets cover the diversity of challenges faced in graph learning and highlights the importance of evaluating models across multiple types of distribution shifts.
7 Future Directions
Theoretical Understanding of Graph Distribution Shifts. Future theoretical analyses could delve deeper into the feasibility and effectiveness of graph OOD adaptation, particularly in scenarios where the label function changes between training and test data. There is also a need to develop theories and methodologies specifically tailored for graph data or graph models, taking into account the intricate structural information inherent in graphs [172]. Furthermore, it is worth exploring more diverse scenarios, such as universal domain adaptive node classification [173], graph size adaptation [27], and multi-source transfer. Notably, several generalization bounds are derived from the graph transferability evaluation perspective [174], [6], [175], and may assist in selecting high-quality source graphs in the multi-source transfer setting. However, identifying the optimal combination of source graphs with theoretical guarantees remains an open problem.
Label-efficient Test-time Graph OOD Adaptation. Test-time adaptation has garnered increasing attention in traditional machine learning, yet relatively few works have been conducted for graph settings. One major challenge is performing test-time adaptation effectively with limited labeled data. When only a few labeled instances are available, there can be significant distribution shifts between the labeled data and and the vast amount of unlabeled test data. With sparse labeled examples, the adaptation process may struggle to accurately reflect the underlying distribution of the target domain. At the same time, if the labeled data do not adequately represent the test distribution, the model’s ability to generalize effectively from these limited labels is compromised. To address these challenges, future research could explore techniques such as semi-supervised learning, active learning, and self-training approaches specifically tailored for graph data.
Trustworthy Graph Learning under Distribution Shifts. Although research in trustworthy graph learning (e.g., fairness [176], explainability [156], reliability [177]) has been extensively studied in recent years, tackling distribution shifts in trustworthy graph learning remains under-explored. Current notations, methodologies, and metrics on trustworthiness may not be directly applicable to scenarios with distribution shifts. For instance, a fair graph learning model might not achieve the same level of fairness when generalizing it to an unseen graph. This calls for the development of new theories and models specifically designed for trustworthy graph learning in OOD and adaptation scenarios. Addressing these gaps could significantly advance reliable graph-based decision-making in dynamic and shifting environments.
Distribution Shifts on Complex Graphs. In contrast to the substantial efforts dedicated to addressing distribution shifts on regular graphs, studies on more complex graph types, such as spatial, temporal, spatial-temporal, heterogeneous graphs, have received comparatively less attention [178]. Such complex graphs often exhibit diverse and dynamic patterns or involve entities and relationships of various types, introducing more intricate and nuanced distribution shifts. The comprehensive exploration and efficient mitigation of distribution shifts on complex graph types are pivotal for enhancing the capabilities of graph machine learning in broader scenarios, such as recommendation systems, healthcare systems, and traffic forecasting.
Graph Foundation Model for Combating Distribution Shift. Developing foundation models specifically tailored for graphs holds significant potential for addressing distribution shifts. These models could integrate comprehensive graph representation learning with adaptation strategies to enhance generalization across varying domains [179]. From the algorithm perspective, developing graph foundation models involves creating architectures that are flexible and adaptive to distribution shifts with minimal fine-tuning. This allows for efficient adaptation to new distributions while preserving the integrity of learned representations. Furthermore, efficient fine-tuning methods, such as graph prompt tuning, are essential for maintaining the generalization capabilities of these models in dynamic, real-world applications [180]. Such techniques can ensure that the models remain adaptable and effective when faced with shifts in graph topology and attributes.
8 Conclusion
In this survey, we comprehensively analyze existing methods related to deep graph learning under distribution shifts, covering three critical scenarios: graph OOD generalization, training-time graph OOD adaptation, and test-time graph OOD adaptation. First, we establish problem definitions and explore different graph distribution shift types that can occur. Following this, we discuss related topicsand explain our categorization. Based on the proposed taxonomy, we systematically review techniques designed to enhance model generalizability or mitigate distribution shifts, exploring both model-based and data-centric approaches. Furthermore, we provide a summary of the datasets commonly utilized in this research area, highlighting their significance and relevance to the challenges associated with distribution shifts. Finally, we emphasize several challenges that remain and outline promising future research directions. We hope that this survey will serve as a valuable resource for researchers seeking to deepen their understanding of the current progress and trends in graph OOD learning.